OpenStudy (anonymous):

If x is nine less than seven and y is nine less seven, what is x less y? Please show me the steps.

OpenStudy (anonymous):

\[x=7-9=-2\]\[y=7-9=-2\]\[y-x=-2-(-2)=-2+2=0\]

OpenStudy (anonymous):

x=7-9;y=7-9 y-x=(7-9)-(7-9) y-x=2-2 so, y-x=0

OpenStudy (anonymous):

1)"If x is nine less than seven" i.e x=7-9 which makes x= -2 2)"y is nine less seven" i.e y=7-9 which makes y= -2 3) "what is x less y?" i.e y-x=? (now substitute the values of x= -2 and y= -2 ) therefore, -2+2= 0